# A broken debris cascade as a possible source of hot dust emission in transitioning planet-forming disks

### v1

@@ Abstract
Context. Planet-forming disks turn from gas-rich, massive disks made of dust and gas into planetary systems containing only small amounts dust produced by collisions between smaller planetary objects like planetesimals, asteroids, or comets. Traditionally we talk about protoplanetary (age $\sim 1$ Myr), transitional ($\sim$ 5–10 Myr), and debris disks ($\sim$ 10-hundreds of Myr), even though the overlap between these phases may be relevant.
Aims. We aim to show that in the transition phase of a disk, when gas surface densities are reduced but not yet negligible, a seemingly small amount of collisional activity may lead to the production of dust on a level that is observationally relevant by creating regions of optical depth 1 or above. In particular, we aim to show that the hot dust emission component of transitional disks may in fact be debris dust produced in such collisions.
Methods. We develop an analytical model to derive the conditions under which observationally relevant amounts of dust can be produced. We focus on the effect of the gas surface density during the transition phase of a disk from fully gas-dominated to the gas-poor debris stage.
Results. We show that the decrease of the gas surface density has an important effect. It allows smaller planetesimals to become collisional, starting a cascade. At the same time, small particles are not destroyed by collisions. That interrupted cascade is critical to preserve the produced dust for significant time intervals, allowing a seemingly minor amount of planetesimal collisions to be effective in producing detectable amounts of dust.
Conclusions. The warm emission of low amounts of dust in many transitional planet-forming disks might be caused entirely by second-generation dust, exposing planetary material at a time much earlier than the age of what are traditionally considered as debris disks.
@@ 1 Introduction
The classical picture of the formation of a planetary systems divides the process into a number of steps: (i) collapse of a molecular core in a molecular cloud, (ii) disk formation due to conservation of angular momentum during collapse as well as due to viscous spreading, (iii) planet formation in the disk by core-accretion and possibly gravitational instabilities, (iv) clearing out the gas in the disk by a combination of viscous accretion onto the star, planet formation and photoevaporative and/or magnetically driven outflows (winds), and (v) a mature planetary system in which rings of planetesimals can produce dust in collisions, ushering in the debris disk phase of the disk that can last for gigayears.
@@
One of the puzzling observations in young planet-forming disks is the observation of a significant component of near infrared emission that is attributed to very warm dust near typical sublimation temperatures around 1500 K. That emission is indeed located close to the star (e.g. Benisty et al., 2011). That excess emission was originally interpreted as being due to a disk inner region heated by accretion (Hillenbrand et al., 1992; D’Alessio et al., 1998, 1999). Natta et al. (2001a) and Dominik et al. (2003a) realised, that the inner edge of a disk is much more intensely irradiated by that star than a disk surface, leading to enhanced emission also in disks that are heated solely by stellar irradiation. Even though a full recovery of the measured near-IR emission requires enhanced scale heights in the inner disk, the basic assumption has been that these processes are responsible for the emission seen.
@@
Meeus et al. (2001) established that there are two basic types of Spectral Energy Distributions (SED) for young Herbig disks that are distinguished by the amount of mid-IR emission, but that show quite similar properties in the near-IR. The assumption was that the these disks differ in the amount of flaring (group-I versus group II disks, Dullemond & Dominik (2004)), but that the properties of the innermost edge of the disk at the sublimation point are very similar, with the infrared emission amounting to 10-30% of the stellar flux (Dominik et al., 2003b). However, eventually it became clear that the group-I disk SED is a symptom of a large inner hole in the disk (Maaskant et al., 2013), a hole like one would expect to be carved out by a forming planetary system. The basic picture is then that a massive planet creates a dust trap in a pressure bump outside the planetary orbit, keeping most of the dust from flowing into the inner disk region. Gas can still pass the planetary orbit through streamers, but the flow of dust should be substantially reduced. This then throws some doubt on the interpretation of the near-IR excess. Why is it then that there is a nearly empty cavity, but at the inner edge, close to the star, the dust suddenly re-appears? In viscous accretion disks, the accretion timescales are shortest in the inner disk, and one would expect the hole to be largely emptied out.
@@
A number of ideas have been put forward to explain the detected emission even in the presence of a strongly dust-depleted gap. Casassus et al. (2013) detected fast streamers in the cavity of HD142527. These streamers can be created through the interaction with the observed companion of that star (Price et al., 2018). If these streamers are narrow and fast, this may account for the low emission coming from the region between inner and outer disk. Pinilla et al. (2016) put forward the idea that small dust grains may sneak through the trap, quickly grow again to larger sizes that are harder to detect, and then fall apart at the snowline. There the water ice that may have been binding dust aggregates evaporates, leading to the sudden appearance of small grains that drift more slowly. The resulting traffic jam is then responsible for the observed near-IR emission.
@@
In this paper, we will explore the idea that the dust seen at the inner edge of these transition disks is not primary dust funneled from the dust trap into the inner system, but that it is secondary dust, produced by the collisions of planetesimals. This would give these disks at least partially the status of debris disk.
@@
There is a large body of literature on dust production in debris disks. In addition to a number of analytical estimates and computations (e.g. Wyatt & Dent, 2002; Dominik & Decin, 2003; Krivov et al., 2006a; Wyatt et al., 2007b, a; Löhne et al., 2008; Kobayashi & Tanaka, 2010; Kennedy & Wyatt, 2010; Wyatt et al., 2011), numerical modeling of debris disks over long times is an established field (e.g. Thébault et al., 2003; Krivov et al., 2005, 2006b; Thébault & Augereau, 2007; van Lieshout et al., 2014). Other papers put the production of debris into the context of a model starting with a gas-rich disk, focusing on the effect of the gas on the larger parent bodies for dust production (e.g. Kenyon & Bromley, 2005, 2008, 2010; Raymond et al., 2011, 2012; Kobayashi & Löhne, 2014). The onset of debris formation in gas-rich disks during planet formation has been proposed in a set of two papers focusing on the disk of HD 163296 (Turrini et al., 2019; D’Angelo & Marzari, 2022), and studied further in a parameter study (Bernabò et al., 2022). These are very detailed models that consider the scattering of planetesimals by the forming giant planets in the system, and they estimate the amount of dust production resulting from these collisions, zooming in on a specific disk.
@@
In the present paper, we are asking a more focused, basic question, namely what conditions are needed in order to create a torus of moderate optical depth in the inner regions of a disk, such as it would be needed to reproduce a low-mass, high scale-height inner disk. We do this by developing an analytical model that captures the dust production of a cascade in the presence of gas, considering that the gas may stop the collisional cascade for the smallest particles. In particular, we discuss the relevance of a reduced gas surface density to constrain when such a broken cascade can occur and what the relation between the mass input from collisions and the optical depth of the torus are.
@@
In Sect. 2 we show the effect of drag on the collisional cascade, derive the important timescales associated with it and show how it breaks the cascade. These effects are applied in Sect. 3 to derive the size distribution of particles in the system. In Sect. 4 we calculate the amount of particles needed to create a moderately optically thick cloud of dust, and along with it show how much mass flux of incoming planetesimals is needed to sustain such a structure. In Sect. 5 we evaluate the derived equations, derive and discuss the outcome and put it into context. Finally, in Sect. 6 we summarize our conclusions.
2 The effect of gas on the cascade
@@ 2.1 Introduction
Figure 1: The proposed model consists of a dust cloud located a distance $R$ from the star (with a width $\Delta R$), suspended in a gaseous medium. A planet is scattering planetesimals out of the midplane, which collide and fragment. The smallest particles are bound to the gas and slowly settle into the midplane.
@@
Fig. 1 shows the basic model: close in to the star, a cloud of dust suspended in the disk gas covers a large solid angle $\alpha$ around the star. We assume that we are dealing with a torus-like structure that contains a mix of gas, dust, and planetesimals. A planet scatters planetesimals into highly inclined orbits, which start to collide and fragment into smaller and smaller particles. Particles below a certain size $s_{\mathrm{b}}$ are tightly bound (coupled) to the gas and start to move along with it, removing relative velocities between them, preventing them from further fragmentation. Instead, these particles slowly settle into the midplane. Once located in the midplane, we assume that these particles no longer contribute to the system. midplane particles are basically irrelevant for the amount of stellar radiation reprocessing, because the maximum amount of stellar radiation they could absorb even under optimal conditions (optically thick geometrically thin midplane of dust in an otherwise optically thin disk) becomes extremely small unless the disk extends all the way to the star. For an inner disk radius of 0.1 au, that fraction is about 2%. For an inner radius of 1 au (close to the sublimation radius in the disks of Herbig stars), only $10^{-3}$ of the stellar photons are absorbed in the midplane (Natta et al., 2001b). These numbers are too small compared to the 10s of percent we are looking for. We also assume that these particles become irrelevant for the collisional cascade as they may re-coagulate and drift into the sublimation zone.
@@
Figure 2: A schematic log-log representation of the timescales governing particles of a certain size (upper part of the plot) and of the resulting abundances (lower part) of particles at a given size. These abundances are integrated over the torus, but excluding particles that have fully settled to the midplane and are treated as irrelevant. The collisional lifetime (yellow line) of a particle increases with size. Particles smaller than $s_{\mathrm{b}}$ are tightly bound to the gas, and are dragged into the midplane by gravity on a settling timescale (light blue) which decreases with particle size. For larger particles, the inclination of their orbit is damped on a damping timescale (dark blue), which increases with particle size. Particles smaller than $s_{\mathrm{cc}}$ have a damping timescale shorter than their collisional timescale, removing them from the collisional cascade. The effect of increasing gas density on these timescales is shown by the dashed lines.
@@
Whereas the smallest particles below $s_{\mathrm{b}}$ couple to the gas and completely leave their inclined orbits around the star, larger particles retain their orbit. The effect of gas does however slowly dampen their inclination, moving them into the midplane as well. There is some range in the particle sizes which do not couple to the gas, but for which damping occurs quicker than collisions, essentially removing them from the cascade as well. The size above which the collisional timescale is lower, and particles are in the collisional cascade we call $s_{\mathrm{cc}}$. These timescales and boundary sizes are shown in Fig. 2 and are derived below.
@@
2.2 Coupling, settling & damping
In order to derive what particles are bound to the gas, we consider Epstein drag, given by Armitage (2010) as
@@ 2.2 Coupling, settling & damping
with $\rho_{\mathrm{gas}}$ the gas density, $s$ the particle size, $v$ the particle velocity relative to the gas and the thermal velocity of gas molecules (Armitage, 2010)
@@
The timescale over which the gas significantly affects a particle’s motion, the friction-time, is given by Armitage (2010) as
@@
Particles are considered tightly bound to the gas if their Stokes number $\tau_{\mathrm{fric}}$, defined as the friction time times the local Kepler frequency, is smaller than 1/2 (see appendix B for the origin of this definition),
$\tau_{\mathrm{fric}}\equiv t_{\mathrm{fric}}\Omega_{\mathrm{K}}<1/2$ (6)
@@
While mainly moving along with the gas, these small particles are being pulled down by gravity, causing them to settle into the midplane. According to Armitage (2010), the settling occurs at a settling velocity found from equating the vertical component of gravity to the drag from this vertical motion, i.e.
@@
Using Eq. (4) for the mass of a particle and substituting the Kepler frequency
@@
we find for the settling velocity
@@
Using this, the timescale at which a particle settles into the midplane becomes
@@ 2.3 Collisional lifetime
As described by Dominik & Decin (2003), the collisional lifetime $t_{\mathrm{coll}}$ of a single particle in the collisional cascade is given by the volume $V$ in which the particles are located, divided by the volume sweeped by the particle per unit time
@@
with $\nu v_{\mathrm{K}}$ the relative velocity between particles in terms of the Kepler velocity $v_{\mathrm{K}}$ and $\sigma_{\mathrm{tot}}$ the total collisional cross-section. Below we will be making the implicit assumption that the relative velocities result from inclinations of the orbits. In reality, a similar component of relative velocities stems from eccentricities of orbits. However, the excitation and damping factors affect both inclinations and eccentricities in similar ways, so we will assume that the inclinations are a good proxy for the source of relative velocities. If the smallest particle capable of destroying a particle of size $s$ has size $\epsilon s$, the total collisional cross-section of a particle is given by
@@
with $f_{\mathrm{cc}}(s)$ the size distribution of particles in a steady-state collisional cascade and $s_{\mathrm{pt}}$ the size of the largest planetesimal in the cascade. $\epsilon$ is given by Dominik & Decin (2003) as
$\epsilon^{3}=\frac{(\nu v_{\mathrm{K}})^{2}}{4S}-\sqrt{\frac{(\nu v_{\mathrm{K}})^{4}}{16S^{2}}-\frac{(\nu v_{\mathrm{K}})^{2}}{2S}}-1$ (16)
@@
with $f_{0,\mathrm{cc}}$ a constant and $\gamma_{\mathrm{cc}}=-3.5$. Since $\gamma_{\mathrm{cc}}<-3$ the cross-section is dominated by the smallest grains, and the integral above evaluates to
@@
Now, by substitution into Eq. (20) and using $\gamma_{\mathrm{cc}}=-3.5$ we find the collisional lifetime of a particle to be

### v3

@@ Abstract
Context. Planet-forming disks go from gas-rich, massive disks made of dust and gas into planetary systems containing only small amounts dust. This dust is produced by collisions between smaller planetary objects, such as planetesimals, asteroids, and comets. Traditionally, we talk about protoplanetary (age $\sim 1$ Myr), transitional ($\sim$ 5–10 Myr), and debris disks ($\sim$ 10-hundreds of Myr) even though the overlap between these phases may be relevant.
Aims. We aim to show that in the transition phase of a disk, when the gas surface densities are reduced but not yet negligible, a seemingly small amount of collisional activity may lead to the production of dust on a level that is observationally relevant by creating regions characterised by an optical depth of 1 or above. In particular, we aim to show that the hot dust emission component of transitional disks may in fact be debris dust that has been produced in such collisions.
Methods. We developed an analytical model to derive the conditions in which observationally relevant amounts of dust can be produced. We focussed on the effect of the gas surface density during the transition phase of a disk from fully gas-dominated to the gas-poor debris stage.
Results. We show that the decrease in the gas surface density has an important effect. It allows for smaller planetesimals to become collisional, initiating a cascade. At the same time, small particles are not destroyed by collisions. This interrupted cascade is critical in terms of preserving the produced dust for significant time intervals, allowing for a seemingly minor amount of planetesimal collisions to be effective in producing detectable amounts of dust.
Conclusions. The warm emission of low amounts of dust in many transitional planet-forming disks might be caused entirely by second-generation dust, exposing planetary material at a much earlier time than the age that debris disks have traditionally been characterised by.
@@ 1 Introduction
The classical picture of the formation of a planetary systems divides the process into a number of steps: (i) the collapse of a molecular core in a molecular cloud, (ii) disk formation due to conservation of angular momentum during collapse as well as due to viscous spreading, (iii) planet formation in the disk by core-accretion and possibly gravitational instabilities, (iv) clearing out the gas in the disk by a combination of viscous accretion onto the star, planet formation, and photoevaporative and/or magnetically driven outflows (winds), and (v) mature planetary system in which rings of planetesimals can produce dust in collisions, ushering in the debris disk phase of the disk, which can last for gigayear scales.
@@
One of the more puzzling observations in young planet-forming disks is the observation of a significant component of near infrared (NIR) emission that is attributed to very warm dust near typical sublimation temperatures around 1500 K. This emission is indeed located close to the star (e.g., Benisty et al., 2011). It was originally interpreted as being due to a disk inner region heated by accretion (Hillenbrand et al., 1992; D’Alessio et al., 1998, 1999). However, Natta et al. (2001a) and Dominik et al. (2003a) realised, that the inner edge of a disk is much more intensely irradiated by that star than a disk surface, leading to enhanced emission also in disks that are heated solely by stellar irradiation. Even though a full recovery of the measured NIR emission requires enhanced scale heights in the inner disk, the basic assumption has been that these processes are responsible for the observed emission.
@@
Meeus et al. (2001) established that there are two basic types of spectral energy distributions (SED) for young Herbig disks that are distinguished by the amount of mid-IR emission, but that also show quite similar properties in the NIR. The assumption stated that these disks differ in the amount of flaring (group-I versus group II disks, Dullemond & Dominik (2004)), but that the properties of the innermost edge of the disk at the sublimation point are very similar, with the IR emission amounting to 10-30% of the stellar flux (Dominik et al., 2003b). However, eventually it became clear that the group-I disk SED is a symptom of a large inner hole in the disk (Maaskant et al., 2013); this hole is akin to the one we would expect to be carved out by a forming planetary system. The basic picture is then that a massive planet creates a dust trap in a pressure bump outside the planetary orbit, keeping most of the dust from flowing into the inner disk region. Gas can still pass the planetary orbit through streamers, but the flow of dust would be substantially reduced. This then throws some doubt on the interpretation of the NIR excess. Consequently, the question arises of why, even though there is a nearly empty cavity, the dust suddenly re-appears at the inner edge, close to the star. In viscous accretion disks, the accretion timescales are shortest in the inner disk and we would expect the hole to be largely emptied out.
@@
A number of hypotheses have been put forward to explain the detected emission, even in the presence of a strongly dust-depleted gap. Casassus et al. (2013) detected fast streamers in the cavity of HD142527. These streamers can be created through the interaction with the observed companion of that star (Price et al., 2018). If these streamers are narrow and fast, this may account for the low emission coming from the region between inner and outer disk. Pinilla et al. (2016) put forward the idea that small dust grains may sneak through the trap, quickly grow again to larger sizes that are harder to detect, and then fall apart at the snowline. There the water ice that may have been binding dust aggregates evaporates, leading to the sudden appearance of small grains that drift more slowly. The resulting traffic jam is then responsible for the observed NIR emission.
@@
In this paper, we explore the idea that the dust seen at the inner edge of these transition disks is not primary dust funneled from the dust trap into the inner system, but that it is secondary dust, produced by the collisions of planetesimals. This would give these disks (at least partially) the status of debris disks.
@@
There is a large body of literature on dust production in debris disks. In addition to a number of analytical estimates and computations (e.g., Wyatt & Dent, 2002; Dominik & Decin, 2003; Krivov et al., 2006a; Wyatt et al., 2007b, a; Löhne et al., 2008; Kobayashi & Tanaka, 2010; Kennedy & Wyatt, 2010; Wyatt et al., 2011), numerical modeling of debris disks over long times is an established field (e.g., Thébault et al., 2003; Krivov et al., 2005, 2006b; Thébault & Augereau, 2007; van Lieshout et al., 2014). Other works have put the production of debris into the context of a model starting with a gas-rich disk, focussing on the effect of the gas on the larger parent bodies for dust production (e.g., Kenyon & Bromley, 2005, 2008, 2010; Raymond et al., 2011, 2012; Kobayashi & Löhne, 2014). The onset of debris formation in gas-rich disks during planet formation has been proposed in a set of two papers focussed on the disk of HD 163296 (Turrini et al., 2019; D’Angelo & Marzari, 2022) and studied further in a parameter study (Bernabò et al., 2022). These are very detailed models that consider the scattering of planetesimals by the forming giant planets in the system, and they estimate the amount of dust production resulting from these collisions, zooming in on a specific disk.
@@
In the present paper, we are asking a more focussed and basic question, namely: what conditions are needed in order to create a torus of moderate optical depth in the inner regions of a disk required to reproduce a low-mass, high scale-height inner disk. We achieved this by developing an analytical model that captures the dust production of a cascade in the presence of gas, while considering that the gas may stop the collisional cascade for the smallest particles. In particular, we discuss the relevance of a reduced gas surface density to constrain when such a broken cascade can occur and consider the relation between the mass input from collisions and the optical depth of the torus.
@@
In Sect. 2, we show the effect of drag on the collisional cascade, derive the important timescales associated with it, and show how it breaks the cascade. These effects are applied in Sect. 3 to derive the size distribution of particles in the system. In Sect. 4, we calculate the amount of particles needed to create a moderately optically thick cloud of dust; alongside it, we show how much mass flux of incoming planetesimals is needed to sustain such a structure. In Sect. 5, we evaluate the derived equations, derive and discuss the outcome and put it into context. Finally, in Sect. 6, we summarise our conclusions.
2 Effect of gas on the cascade
@@ 2.1 Introduction
Figure 1: Proposed model consists of a dust cloud located a distance, $R,$ from the star (with a width $\Delta R$), suspended in a gaseous medium. A planet scatters planetesimals out of the midplane, which collide and fragment. The smallest particles are bound to the gas and slowly settle into the midplane.
@@
Figure 1 shows the basic model: close in to the star, a cloud of dust suspended in the disk gas covers a large solid angle $\alpha$ around the star. We assume that we are dealing with a torus-like structure that contains a mix of gas, dust, and planetesimals. A planet scatters planetesimals into highly inclined orbits, which start to collide and fragment into smaller and smaller particles. Particles below a certain size, $s_{\mathrm{b}}$, are tightly bound (coupled) to the gas and start to move along with it, removing relative velocities between them, preventing further fragmentation. Instead, these particles slowly settle into the midplane. Once located in the midplane, we assume that these particles no longer contribute to the system. Midplane particles are basically irrelevant for the amount of stellar radiation reprocessing, because the maximum amount of stellar radiation they could absorb even under optimal conditions (optically thick geometrically thin midplane of dust in an otherwise optically thin disk) becomes extremely small unless the disk extends all the way to the star. For an inner disk radius of 0.1 au, that fraction is about 2%. For an inner radius of 1 au (close to the sublimation radius in the disks of Herbig stars), only $10^{-3}$ of the stellar photons are absorbed in the midplane (Natta et al., 2001b). These numbers are too small compared to the tens of percent we are looking for. We also assume that these particles become irrelevant for the collisional cascade as they may re-coagulate and drift into the sublimation zone.
@@
Figure 2: Schematic log-log representation of the timescales governing particles of a certain size (top of the plot) and of the resulting abundances (bottom) of particles at a given size. These abundances are integrated over the torus, but excluding particles that have fully settled to the midplane and are treated as irrelevant. The collisional lifetime (yellow line) of a particle increases with size. Particles smaller than $s_{\mathrm{b}}$ are tightly bound to the gas, and are dragged into the midplane by gravity on a settling timescale (light blue) which decreases with particle size. For larger particles, the inclination of their orbit is damped on a damping timescale (dark blue), which increases with particle size. Particles smaller than $s_{\mathrm{cc}}$ have a damping timescale shorter than their collisional timescale, removing them from the collisional cascade. The effect of increasing gas density on these timescales is shown by the dashed lines.
@@
While the smallest particles below $s_{\mathrm{b}}$ couple to the gas and completely leave their inclined orbits around the star, larger particles retain their orbit. The effect of gas does however slowly dampen their inclination, moving them into the midplane as well. There is some range in the particle sizes which do not couple to the gas, but for which damping occurs quicker than collisions, essentially removing them from the cascade as well. The size above which the collisional timescale is lower, and particles are in the collisional cascade we refer to as $s_{\mathrm{cc}}$. These timescales and boundary sizes are shown in Fig. 2 and their derivation is described below.
@@
2.2 Coupling, settling, and damping
To derive what particles are bound to the gas, we consider the Epstein drag, given by Armitage (2010) as:
@@ 2.2 Coupling, settling, and damping
with $\rho_{\mathrm{gas}}$ the gas density, $s$ the particle size, $v$ the particle velocity relative to the gas, and the thermal velocity of gas molecules (Armitage, 2010) given as:
@@
The timescale over which the gas significantly affects a particle’s motion, the friction-time, is given by Armitage (2010) as:
@@
Particles are considered tightly bound to the gas if their Stokes number $\tau_{\mathrm{fric}}$, defined as the friction time times the local Kepler frequency, is smaller than 1/2 (see Appendix B for the origin of this definition), as follows:
$\tau_{\mathrm{fric}}\equiv t_{\mathrm{fric}}\Omega_{\mathrm{K}}<1/2,$ (6)
@@
Although they are mainly moving along with the gas, these small particles are being pulled down by gravity, causing them to settle into the midplane. According to Armitage (2010), the settling occurs at a settling velocity found from equating the vertical component of gravity to the drag from this vertical motion, namely:
@@
Using Eq. (4) for the mass of a particle and substituting the Kepler frequency:
@@
for the settling velocity, we find:
@@
On this basis, the timescale at which a particle settles into the midplane becomes:
@@ 2.3 Collisional lifetime
As described by Dominik & Decin (2003), the collisional lifetime, $t_{\mathrm{coll}}$, of a single particle in the collisional cascade is given by the volume $V$, in which the particles are located, divided by the volume sweeped by the particle per unit time:
@@
with $\nu v_{\mathrm{K}}$ as the relative velocity between particles in terms of the Kepler velocity, $v_{\mathrm{K}}$, and $\sigma_{\mathrm{tot}}$ the total collisional cross-section. Below, we make the implicit assumption that the relative velocities result from inclinations of the orbits. In reality, a similar component of relative velocities stems from the eccentricities of orbits. However, the excitation and damping factors affect both inclinations and eccentricities in similar ways, so we assume that the inclinations are a good proxy for the source of relative velocities. If the smallest particle capable of destroying a particle of size $s$ has a size of $\epsilon s$, the total collisional cross-section of a particle is given by:
@@
with $f_{\mathrm{cc}}(s)$ the size distribution of particles in a steady-state collisional cascade and $s_{\mathrm{pt}}$ the size of the largest planetesimal in the cascade. Then, $\epsilon$ is given by Dominik & Decin (2003) as:
$\epsilon^{3}=\frac{(\nu v_{\mathrm{K}})^{2}}{4S}-\sqrt{\frac{(\nu v_{\mathrm{K}})^{4}}{16S^{2}}-\frac{(\nu v_{\mathrm{K}})^{2}}{2S}}-1,$ (16)
@@
with $f_{0,\mathrm{cc}}$ a constant and $\gamma_{\mathrm{cc}}=-3.5$. Since $\gamma_{\mathrm{cc}}<-3$ the cross-section is dominated by the smallest grains, and the integral above is evaluated as:
@@
Now, by substituting this into Eq. (20) and using $\gamma_{\mathrm{cc}}=-3.5,$ we find the collisional lifetime of a particle to be: